OpenStudy (anonymous):

evaluate the integral: (5x+1)/((2x+1)(x-1))

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=%285x%2B1%29%2F%28%282x%2B1%29%28x-1%29%29 Click on show steps..

OpenStudy (anonymous):

solve it by a partial fraction method

myininaya (myininaya):

\[\frac{5x+1}{(2x+1)(x-1)}=\frac{A}{2x+1}+\frac{B}{x-1}\] \[5x+1=A(x-1)+B(2x+1)=x(A+2B)+(-A+B)=5x+1\] \[ => A+2B=5, -A+B=1 \] \[-A+B=1=> B=1+A \] \[A+2B=5=> A+2(1+A)=5 => 3A+2=5=> A=1\] \[A=1=> B=1+A=1+1=2\] so we have \[\int\limits_{}^{}\frac{5x+1}{(2x+1)(x-1)} dx=\int\limits_{}^{}(\frac{1}{2x+1} +\frac{2}{x-1}) dx\] \[=\frac{1}{2}\ln|2x-1|+2 \ln|x-1|+C\]

OpenStudy (anonymous):

=>A+2B=5,−A+B=1 −A+B=1=>B=1+A How do u know that?

myininaya (myininaya):

if you want wolfram's answer you need to play with ln|x-1|: ln|x-1|=ln|(-1)(-x+1)|=ln|(-1)(1-x)| =ln[|-1| *|1-x|] =ln[|1|*|1-x|]=ln(1)+ln|1-x|=0+ln|1-x|=ln|1-x|

myininaya (myininaya):

right hand side must = left and side x coefficient on left side must = x coefficient on right side constant term on left side must=constant term on right side

OpenStudy (anonymous):

1/2ln|2x−1|+2ln|x−1|+C where did that first 1/2 come from?

myininaya (myininaya):

\[\int\limits_{}^{}\frac{1}{2x+1} dx\] \[u=2x+1=> du =2 dx => \frac{1}{2} du=dx\] \[\int\limits_{}^{}\frac{1}{2x+1} dx=\int\limits_{}^{}\frac{1}{2} \frac{du}{u}=\frac{1}{2} \ln|u|+C\]

OpenStudy (anonymous):

oh ok, yeah i guess im just still lost on the partial fractions...i dont quite understand how we ended up with 1 and 2

myininaya (myininaya):

we got two equations with 2 unknowns and solved one of the equations for B and plugged into the other equation and solved for A

OpenStudy (anonymous):

oh! ok that makes sense!